# Strong-to-weak spontaneous breaking of 1-form symmetry and intrinsically mixed topological order

### v2

@@ II.3 Strong and weak 1-form symmetries and order/disorder parameters
$D_{1}=\lim\limits_{|i-j|\to\infty}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|U_{g,u}(i,j)|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (26)
@@
$O_{1}=\lim\limits_{r\to\infty}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|U_{h,u}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ (28)
$D_{2}=\lim\limits_{|i-j|\to\infty}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|U_{g,u}(i,j)U_{g,l}^{T}(i,j)|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ (29)
@@ II.4 Emergent 1-form symmetries
$\tilde{U}_{g}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\propto|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ (33)
We define the topological order of the Choi state by the algebraic relations (braiding, fusion, etc) of the emergent 1-form symmetries, for which $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ is an exact eigenstate. The phase of the density matrix then follows from the phase of its Choi state. [^14] We do not assume that the definition of phase for the Choi state matches in this context exactly with the definition for usual pure gapped ground states. This is because the action of finite time Lindbladian evolution preserving finite Markov length on a density matrix, when translated to a unitary action on the Choi state, might not be finite time evolution of a local Hamiltonian. Here, we simply say that two Choi states are in the same “phase” if they can be connected by action corresponding to finite Lindbladian evolution, i.e. a sequence of infinitesimal locally generated non-unitary evolutions appropriately normalized, without a diverging correlation length.
@@ III Phases of mixed states: $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1-form symmetry
$\rho=4\prod\left(\frac{1+B_{p}}{2}\right)\prod\left(\frac{1+A_{v}}{2}\right)$ (35)
@@
The intrinsically mixed fixed point is described by
$\rho=\frac{2^{N_{p}}}{2^{N_{e}}}\prod\left(\frac{1+B_{p}}{2}\right)$ (37)
Computing the order and disorder parameters in this state gives
$\begin{split}&O_{2,x}=1\qquad D_{1,z}=0\\ &O_{1,x}=0\qquad D_{2,z}=1\\ \end{split}$ (38)
@@
[…]
@@
$\rho=\frac{1}{2^{N_{e}}}\mathbf{1}$ (39)
It only has weak Wilson and ’t Hooft symmetries, and is symmetric under both of them. The density matrix has LW CMI 00 and the Choi state has TEE 00. Because the Choi state for the intrinsically mixed phase describes a single copy of toric code and the Choi state for the trivial phase describes a trivial topological order, then there must be a phase transition marked by diverging correlation length in any interpolation between the two states. It follows that the two density matrices are not equivalent via the definition in Sec. II.2; the intrinsically mixed phase is a distinct phase from the trivial phase under this equivalence relation.
@@ IV Example: random vertex term
$\begin{split}&|\psi_{0}(\{\lambda_{A,v}\})\rangle\langle\psi_{0}(\{\lambda_{A,v}\})|\\ &=4\prod_{p}\left(\frac{1+B_{p}}{2}\right)\prod_{v}\left(\frac{1+\mathrm{sign}(\lambda_{A,v})A_{v}}{2}\right)\end{split}$ (41)
The ground state is identical to the ideal toric code ground state, except there is a probability that depends on $\Delta_{A}^{2}$ for a given vertex $v$ to have $\lambda_{A,v}<0$. The ground state then satisfies $A_{v}|\psi_{0}(\{\lambda_{A,v}\})\rangle=-|\psi_{0}(\{\lambda_{A,v}\})\rangle$ for that vertex.
@@
$\rho(\Delta_{A}^{2})=\frac{1}{\cosh^{N}(\beta_{A})}\prod_{p}\left(\frac{1+B_{p}}{2}\right)\prod_{v}e^{\beta_{A}A_{v}}$ (42)
where the effective inverse temperature $\beta_{A}$ is determined by $\Delta_{A}^{2}$. Note that this result depends not only on the amplitude but also the distribution of the disorder; we will show in Sec. IV.1 that if we instead use bimodal disorder, we only obtain density matrices of the form (42) with $\beta=0$ or $\beta=\infty$.
Before we proceed further, it is worth noting that the exact same density matrix can also be found in (1) the toric code at finite temperature with the coefficient $\lambda_{B}$ of the $B_{p}$ term taken to infinity and (2) toric code systems prepared by finite depth circuit, measurement, and error correction (see Refs. Castelnovo and Chamon (2007); Hamma et al. (2009); Iqbal et al. (2024a); Tantivasadakarn et al. (2024); Lee et al. (2022b); Zhu et al. (2023); Iqbal et al. (2023, 2024b)), if there is a nonzero rate of measurement error. In the latter setting, $B_{p}=1$ automatically for all plaquettes due to the choice of measurement basis, as long as the finite-depth circuit has no error. However, $A_{v}=\pm 1$ depending on the outcome of the measurement on the ancillas. One must then correct for the $A_{v}=-1$ vertices by appropriate application of string operators. However, if there is a measurement error where the sign of an ancilla measurement is read off incorrectly, then one generally obtains a toric code state with a small percentage of vertices with $A_{v}=-1$. It follows from the results in this section that for any nonzero rate of measurement error, the resulting density matrix only forms a classical memory, and demonstrates SW-SSB. This gives some physical intuition for why measurement error, even at small probability, is much more detrimental than incoherent noise.
@@ IV.1 Mixed state density matrix and CMI
We begin by deriving the density matrix. Upon averaging over disorder realizations, for any given vertex $v$ there is a probability $\frac{1}{2}\left[1\pm\mathrm{erf}\left(\frac{1}{\sqrt{2}\Delta_{A}}\right)\right]$ to get $\mathrm{sign}(\lambda_{A,v})=\pm 1$. Here, the error function $\mathrm{erf}(x)$ is defined as the integral over the Gaussian distribution from 00 to $x$, normalized so that $\mathrm{lim}_{x\to\infty}\mathrm{erf}(x)=1$. For any given vertex, this probability distribution gives the following density matrix
@@
$\mathrm{tanh}(\beta_{A})=\mathrm{erf}\left(\frac{1}{\sqrt{2}\Delta_{A}}\right).$ (44)
Note that as $\Delta_{A}\to 0,$ $\beta_{A}\to\infty$ returns to the usual toric code model, while as $\Delta_{A}\to\infty,\beta_{A}\to 0$ as expected.
@@
We will now compute the LW CMI of the density matrix (42), which can be carried out exactly because (42) is a product of commuting terms. We define the Rényi-2 LW CMI using the partitions in Fig. 2. To compute the Rényi-2 entanglement entropy between a region and its complement, it is helpful to expand the density matrix in terms of Pauli strings. Using the identity $e^{\beta_{A}O}=\cosh(\beta_{A})+\sinh(\beta_{A})O$ for any operator $O$ that squares to the identity, we obtain
$\rho(\beta_{A})=\frac{1}{2^{N_{e}}}\prod_{p}(1+B_{p})\prod_{v}\left(1+\tanh(\beta_{A})A_{v}\right).$ (45)
If we expand out the product above and trace out a region $\bar{R}$ which is the complement of a simply connected region $R$, then any nontrivial term with support in $\bar{R}$ would be removed by the trace because the trace of every Pauli operator is zero. Therefore, for any simply connected region $R$,
$\begin{split}\rho_{R}(\beta_{A})&=\mathrm{Tr}_{\bar{R}}\left[\rho(\beta_{A})\right]\\ &=\frac{1}{2^{N_{e}}}\prod_{p\in R}\left(1+B_{p}\right)\prod_{v\in R}(1+\tanh(\beta_{A})A_{v}),\end{split}$ (46)
@@ IV.3 Choi state
$\rho_{B}=\frac{1}{2^{N_{e}/2}}\prod_{p}\left(\frac{1+B_{p}}{2}\right),$ (57)
Its Choi state $|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ takes a simple form:
$|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\propto\prod_{e}\left(\frac{1+Z_{e,u}Z_{e,l}}{2}\right)|tc_{u}\rangle|tc_{l}\rangle,$ (58)
We can apply the operator $\prod_{v}e^{-\beta_{A}A_{v}/2}$ on both sides of the density matrix $\rho_{B}$ to obtain our target density matrix
@@
$|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\propto\prod_{v}e^{\beta_{A}A_{v,u}/2}\prod_{v^{\prime}}e^{\beta_{A}A_{v^{\prime},l}/2}|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (60)
@@
We will now work exclusively with spins in $\mathcal{H}_{u}$ and $\mathcal{H}_{m}$. This allows us to rewrite the state $|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ in a simple way:
$|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|tc_{u}\rangle|\uparrow_{m}\rangle,$ (63)
where $|\uparrow_{m}\rangle$ is a product state in the middle Hilbert space $\mathcal{H}_{m}$ with all spins pointing in the $+z$-direction. $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ takes the form
$\begin{split}|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}&\propto\prod e^{\beta_{A}A_{v,u}A_{v,m}/2}\prod e^{\beta_{A}A_{v,m}/2}|tc_{u}\rangle|\uparrow_{m}\rangle\\ &\propto|tc_{u}\rangle\prod e^{\beta_{A}A_{v,m}}|\uparrow_{m}\rangle.\end{split}$ (64)
Therefore, $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ is simply an unperturbed toric code state stacked with a confined state. The state in $\mathcal{H}_{m}$ is confined as easily evaluated from the correlation functions of loops. One can also “ungauge” the $\mathbb{Z}_{2}$ symmetry and find that the state is connected to the ferromagnetic phase for any finite $\beta$. Specifically, the ungauged version of the state in $\mathcal{H}_{m}$ is
@@
The order parameter is $\mathcal{O}(1)$ for all finite $\beta$. Specifically, $\mathrm{Tr}(Z_{i}Z_{j}\rho)=e^{-4\beta_{A}}$, independent of $|i-j|$ for sufficiently separated $i,j$. Furthermore, the disorder parameter evaluates to (dropping the subscript $m$ in the following equation for ease of notation)

### v3

@@ II.3 Strong and weak 1-form symmetries and order/disorder parameters
$D_{1}=\lim\limits_{|i-j|\to\infty}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|U_{g,u}(i,j)|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}},$ (26)
@@
$O_{1}=\lim\limits_{r\to\infty}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|U_{h,u}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (28)
$D_{2}=\lim\limits_{|i-j|\to\infty}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|U_{g,u}(i,j)U_{g,l}^{T}(i,j)|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (29)
@@ II.4 Emergent 1-form symmetries
$\tilde{U}_{g}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\propto|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (33)
We define the topological order of the Choi state by the algebraic relations (braiding, fusion, etc) of the emergent 1-form symmetries, for which $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ is an exact eigenstate. The phase of the density matrix then follows from the phase of its Choi state. [^14] We do not assume that the definition of phase for the Choi state matches in this context exactly with the definition for usual pure gapped ground states. This is because the action of finite time Lindbladian evolution preserving finite Markov length on a density matrix, when translated to a unitary action on the Choi state, might not be finite time evolution of a local Hamiltonian. Here, we simply say that two Choi states are in the same “phase” if they can be connected by action corresponding to finite Lindbladian evolution, i.e. a sequence of infinitesimal locally generated non-unitary evolutions appropriately normalized, without a diverging correlation length.
@@ III Phases of mixed states: $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1-form symmetry
$\rho=\prod_{p}\left(\frac{1+B_{p}}{2}\right)\prod_{v}\left(\frac{1+A_{v}}{2}\right)$ (35)
@@
The intrinsically mixed-state fixed point is described by
$\rho=\frac{1}{2^{N-1}}\prod_{p}\left(\frac{1+B_{p}}{2}\right),$ (37)
where $N$ is the number of vertices. Computing the order and disorder parameters in this state gives
[…]
@@
We note that there is a close relationship between SW-SSB of 0-form symmetry and that of 1-form symmetry. In (2+1)$d$, gauging a 0-form discrete finite global symmetry produces a dual 1-form symmetry generated by the Wilson loop operators of the original symmetry. The above state can be obtained from a density matrix demonstrating SW-SSB of a 0-form non-anomalous $\mathbb{Z}_{2}$ symmetry such as $\rho\propto 1+\prod_{v}X_{v}$ where $X_{v}$ is a Pauli $X$ operator acting on a vertex $v$ (this is simply the 2+1dd version of the 1+1dd example in Ref. Ma et al. (2023)). Eq. (37) can be obtained by gauging the $\mathbb{Z}_{2}$ symmetry $\prod_{v}X_{v}$. This method for constructing 1-form SW-SSB states starting from 0-form SW-SSB states of non-anomalous 0-form symmetry can be generalized to other symmetry groups. See Sec. V.5 for more discussion on the relationship between the topological phases discussed here and Ising spin glass literature.
@@
$\rho=\frac{1}{2^{2N-2}}\mathbf{1}$ (39)
It only has weak Wilson and ’t Hooft symmetries, and is symmetric under both of them. The density matrix has LW CMI 0 and the Choi state has TEE 0. Because the Choi state for the intrinsically mixed phase describes a single copy of toric code and the Choi state for the trivial phase describes a trivial topological order, then there must be a phase transition marked by diverging correlation length in any interpolation between the two states. It follows that the two density matrices are not equivalent via the definition in Sec. II.2; the intrinsically mixed phase is a distinct phase from the trivial phase under this equivalence relation.
@@ IV Example: random vertex term
$\begin{split}&|\psi_{0}(\{\lambda_{A,v}\})\rangle\langle\psi_{0}(\{\lambda_{A,v}\})|\\ &=\prod_{p}\left(\frac{1+B_{p}}{2}\right)\prod_{v\neq v_{0}}\left(\frac{1+\mathrm{sign}(\lambda_{A,v})A_{v}}{2}\right),\end{split}$ (41)
where $v_{0}$ is a fixed vertex where the corresponding value of $A_{v_{0}}$ is determined from the identity $A_{v_{0}}=\prod_{v\neq v_{0}}A_{v}$. The ground state is identical to the ideal toric code ground state, except there is a probability that depends on $\Delta_{A}^{2}$ for a given vertex $v$ to have $\lambda_{A,v}<0$. The ground state then satisfies $A_{v}|\psi_{0}(\{\lambda_{A,v}\})\rangle=-|\psi_{0}(\{\lambda_{A,v}\})\rangle$ for that vertex.
@@
$\rho(\Delta_{A}^{2})=\frac{1}{[2\cosh(\beta_{A})]^{N-1}}\prod_{p}\left(\frac{1+B_{p}}{2}\right)\prod_{v\neq v_{0}}e^{\beta_{A}A_{v}}$ (42)
where the effective inverse temperature $\beta_{A}$ is determined by $\Delta_{A}^{2}$, and $N$ is the number of vertices. Note that this result depends not only on the amplitude but also the distribution of the disorder; we will show in Sec. IV.1 that if we instead use bimodal disorder, we only obtain density matrices of the form (42) with $\beta=0$ or $\beta=\infty$.
Before we proceed further, it is worth noting that the same density matrix can also be found in (1) the toric code at finite temperature with the coefficient $\lambda_{B}$ of the $B_{p}$ term taken to infinity and (2) toric code systems prepared by finite depth circuit, measurement, and error correction (see Refs. Castelnovo and Chamon (2007); Hamma et al. (2009); Iqbal et al. (2024a); Tantivasadakarn et al. (2024); Lee et al. (2022b); Zhu et al. (2023); Iqbal et al. (2023, 2024b)), if there is a nonzero rate of measurement error. In the latter setting, $B_{p}=1$ automatically for all plaquettes due to the choice of measurement basis, as long as the finite-depth circuit has no error. However, $A_{v}=\pm 1$ depending on the outcome of the measurement on the ancillas. One must then correct for the $A_{v}=-1$ vertices by appropriate application of string operators. However, if there is a measurement error where the sign of an ancilla measurement is read off incorrectly, then one generally obtains a toric code state with a small percentage of vertices with $A_{v}=-1$. It follows from the results in this section that for any nonzero rate of measurement error, the resulting density matrix only forms a classical memory, and demonstrates SW-SSB. This gives some physical intuition for why measurement error, even at small probability, is much more detrimental than incoherent noise.
@@ IV.1 Mixed state density matrix and CMI
We begin by deriving the density matrix. Upon averaging over disorder realizations, for any given vertex $v\neq v_{0}$ there is a probability $\frac{1}{2}\left[1\pm\mathrm{erf}\left(\frac{1}{\sqrt{2}\Delta_{A}}\right)\right]$ to get $\mathrm{sign}(\lambda_{A,v})=\pm 1$. Here, the error function $\mathrm{erf}(x)$ is defined as the integral over the Gaussian distribution from 0 to $x$, normalized so that $\mathrm{lim}_{x\to\infty}\mathrm{erf}(x)=1$. For any given vertex, this probability distribution gives the following density matrix
@@
$\mathrm{tanh}(\beta_{A})=\mathrm{erf}\left(\frac{1}{\sqrt{2}\Delta_{A}}\right),$ (44)
where have used the identity $e^{\beta_{A}O}=\cosh(\beta_{A})+\sinh(\beta_{A})O$ for any operator $O$ that squares to the identity. Note that as $\Delta_{A}\to 0,$ $\beta_{A}\to\infty$ returns to the usual toric code model, while as $\Delta_{A}\to\infty,\beta_{A}\to 0$ as expected.
@@
We will now compute the LW CMI of the density matrix (42), which can be carried out exactly because (42) is a product of commuting terms. We define the Rényi-2 LW CMI using the partitions in Fig. 2. To compute the Rényi-2 entanglement entropy between a region and its complement, it is helpful to expand the density matrix in terms of Pauli strings. We have
$\rho(\beta_{A})=\frac{1}{2^{2N-1}}\prod_{p}(1+B_{p})\prod_{v\neq v_{0}}\left(1+\tanh(\beta_{A})A_{v}\right).$ (45)
If we expand out the product above and trace out a region $\bar{R}$ which is the complement of a simply connected region $R$, then any nontrivial term with support in $\bar{R}$ would be removed by the trace because the trace of every Pauli operator is zero. Therefore, for any simply connected region $R$ which, without loss of generality, we assume to not contain $v_{0}$,
$\begin{split}\rho_{R}(\beta_{A})&=\mathrm{Tr}_{\bar{R}}\left[\rho(\beta_{A})\right]\\ &=\frac{1}{2^{2N_{e}}}\prod_{p\in R}\left(1+B_{p}\right)\prod_{v\in R}(1+\tanh(\beta_{A})A_{v}),\end{split}$ (46)
@@ IV.3 Choi state
$\rho_{B}=\frac{1}{2^{N-1}}\prod_{p}\left(\frac{1+B_{p}}{2}\right),$ (57)
Its Choi state $|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ takes a simple form:
$|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\propto\prod_{e}\left(\frac{1+Z_{e,u}Z_{e,l}}{2}\right)|tc_{u}\rangle|tc_{l}\rangle,$ (58)
We can apply the operator $\prod_{v}e^{-\beta_{A}A_{v}/2}$ on both sides of the density matrix $\rho_{B}$ to obtain our target density matrix (the correction due to exclusion of vertex $v_{0}$ will not affect the scaling behavior and will be dropped below)
@@
$|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\propto\prod_{v}e^{\beta_{A}A_{v,u}/2}\prod_{v^{\prime}}e^{\beta_{A}A_{v^{\prime},l}/2}|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}.$ (60)
@@
We will now work exclusively with spins in $\mathcal{H}_{u}$ and $\mathcal{H}_{m}$. This allows us to rewrite the state $|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ in a simple way:
$|\rho_{B}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|tc_{u}\rangle|\uparrow_{m}\rangle,$ (63)
where $|\uparrow_{m}\rangle$ is a product state in the middle Hilbert space $\mathcal{H}_{m}$ with all spins pointing in the +z+z-direction. $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ takes the form
$\begin{split}|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}&\propto\prod e^{\beta_{A}A_{v,u}A_{v,m}/2}\prod e^{\beta_{A}A_{v,m}/2}|tc_{u}\rangle|\uparrow_{m}\rangle\\ &\propto|tc_{u}\rangle\prod e^{\beta_{A}A_{v,m}}|\uparrow_{m}\rangle.\end{split}$ (64)
Therefore, $|\rho(\beta_{A})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ is simply an unperturbed toric code state stacked with a confined state. The state in $\mathcal{H}_{m}$ is confined as easily evaluated from the correlation functions of loops. One can also “ungauge” the $\mathbb{Z}_{2}$ symmetry and find that the state is connected to the ferromagnetic phase for any finite $\beta$. Specifically, the ungauged version of the state in $\mathcal{H}_{m}$ is
@@
The order parameter is $\mathcal{O}(1)$ for all finite $\beta$. Specifically, $\mathrm{Tr}(Z_{i}Z_{j}\rho)=e^{-4\beta_{A}}$, independent of |ij||i-j| for sufficiently separated $i,j$. Furthermore, the disorder parameter evaluates to (dropping the subscript $m$ in the following equation for ease of notation)